Now suppose you encounter three alien species with different genetic material from living things on Earth. The linear relationships between their genetic material composition and melting point can be expressed by the following equations.
y = 3x + 25
y = -6x + 12
y = 1.5x + 17
Drag the correct equation to match each description. The equations may be used once, more than once, or not at all.


1. This equation represents the steepest line.
2. This equation represents a line that slopes down to the right.
3. Of the two equations that slope up to the right, this equation would cross the y-axis closer to the origin.
4. When x = 4 in this equation, y = 37.
1
Expert's answer
2020-06-04T11:17:52-0400

The linear equation can be represented as following:

y = ax + b

where a indicates slope and b is a distance from the origin.

From here:

1. This equation represents the steepest line:

y = -6x + 12. The coefficient a = -6 gives the steepest slope.

2. This equation represents a line that slopes down to the right:

y = -6x + 12. As the coefficient a = -6 is negative, the line slopes down to the right.

3. Of the two equations that slope up to the right, this equation would cross the y-axis closer to the origin:

y = 1.5x + 17. As the coefficient b = 17, the line crosses the y-axis closer to teh origin.

4. When x = 4 in this equation, y = 37.

y = 3x + 25. If x = 4 then 3 × 4 + 25 = 37. y = 37.